A rectangular park having lenght of 72 m and breadth of 54 m is surrounded by a path of 2.5 m wide. Find the area of the path?
Answer by checkley77(12569)

AREA=(72+5)(54+5)-72*54
AREA=77*59-3,888
AREA=4,543-3,888
AREA=655 M^2 IS THE AREA OF THE PATH.